OpenStudy (anonymous):

If f(3) = 14 and f '(x) ≥ 3 for 3 ≤ x ≤ 7, how small can f(7) possibly be?

OpenStudy (anonymous):

@jim_thompson5910 hey jim u think u could help me

myininaya (myininaya):

Looks like we have that mean-value theorem here.

OpenStudy (anonymous):

can I show u what i did? idk where i messed up

myininaya (myininaya):

\[f'(c)=\frac{f(b)-f(a)}{b-a}\] where c is between a and b. (well they are using some number x is in between a and b; just replace c with x here if it makes more since to you)

myininaya (myininaya):

sure

OpenStudy (anonymous):

[f(7)-f(3)/(7-14) [f(7)-14/-7 f(7)-14/-7>=3 c=-7

myininaya (myininaya):

oh no... ok the change of x is 7-3 which is 4. the change of y is f(7)-f(3) and since f(3)=14 we could say f(7)-14

myininaya (myininaya):

so we have that f'(c)>=3 and f'(c)=(f(b)-f(a))/(b-a) so that means (f(b)-f(a))/(b-a)>=3.

myininaya (myininaya):

Your objective is to solve this inequality for f(7)

OpenStudy (anonymous):

ok sec let me try to work it out

OpenStudy (anonymous):

f(7)-14/4>=3

OpenStudy (anonymous):

f(7)=26?

myininaya (myininaya):

almost you had an inequality

myininaya (myininaya):

f(7)>=26

myininaya (myininaya):

This means f(7) is possibly 26 or bigger.

OpenStudy (anonymous):

yay thank you soo much for ur help!

myininaya (myininaya):

Np. You did most of the work. :)